OpenStudy (highschoolmom2010):

Which of these orbital designations are invalid? a)4s b)2d c)3f d)3p

OpenStudy (gebooors):

When n= 2, there are s and p-orbitals When n = 3, there are s, p and d-orbitals 4f is possible

OpenStudy (highschoolmom2010):

i dont understand how you can tell

OpenStudy (ankit042):

read this carefully When n = 3, there are s, p and d-orbitals

OpenStudy (highschoolmom2010):

i still dont understand how to use the model to understand this

OpenStudy (gebooors):

In Bohr´s model, electrons are in energy levels are labeled as K,L,MN,.. MAin quantum number n = 1,2,3,4,.. In quantum mechanical model there are suborbitals s,p,d,f Shape of suborbital is described by side quantum number l. It can have values 0,1,..n-1 for s , l = 0 , for p l = 1 , or d l = 2 and for f l = 3 So, if n = 1 (K) l = 0 (remeber highest value can be n-1) in L, n = 2 and l can be 0 (s) or 1 (p) in M n = 3, and l = 0, 1, 2. 2 is for d suborbital Then N (n = 4) l = 0,1,2,3. For f l = 3 When inner electron levels are filled (each suborbital may have two electrons with opposite spins), there are more suborbitals in outer electron levels.